OpenStudy (kewlgeek555):

9x + 18 = 28x – 4 Steps Please.

OpenStudy (mathstudent55):

Why did you post the same problem again when it is being answered in your earlier post?

OpenStudy (kewlgeek555):

Sorry - you didnt reply in that post until after making this one. I just really need to know how to do this. I have an exam tomorrow.

OpenStudy (anonymous):

I just showed you how to do it but you deleted it

OpenStudy (anonymous):

9x + 18 = 28x – 4 9x + 18 - 28x = 28x -28x -4 -19x + 18 = -4 -19x + 18 - 18 = -4 -18 -19x = -22 -19x / -19 = -22 / -19 x = 22/19

OpenStudy (anonymous):

9x + 18 = 28x – 4 19x+18=-4 19x=-22 x=22/19

OpenStudy (anonymous):

9x + 18 = 28x – 4 37x +18= -4 37x = 14 x =2.64